Yahaira:

 If you add​ Natalie's age and​ Fred's age, the result is 36. If you add​ Fred's age to 3 times​ Natalie's age, the result is 60. Write and solve a system of equations to find how old Fred and Natalie are.

Yahaira:

@mxddi3 pls help

Yahaira:

@XioGonz

XxAngelGirlxX:

well do you know how to solve this problem?

Yahaira:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XxAngelGirlxX well do you know how to solve this problem? \(\color{#0cbb34}{\text{End of Quote}}\) no

mxddi3:

ok so first, this shouldn't be in writing. Anyways, this can be written as 2 equations: X=Natalie y=Fred Now, we write the equations. If their aged added is equal to 36, then `x+y=36` ok then, adding Fred's age to 3 times Natalie's age we get `3x+y=60` so we can use the substitution method for this. So I will start you off, but i need you to finish. So we are going to change `x+y=36` and subtract y from both sides and get `x=36-y` Plug this into the second equation and we get: 3(36-y)+y=60. This simplifies to 108-3y+y=60 That then simplifies to 108-2y=60. Then you need to subtract 108 from both sides and we are left with: `-2y=-48`. Now, you have -2y= -48 so please solve that. You will get your "y" value. Then, from there you take that answer and plug it into the first equation (x+y=36) and you will get your "x" value as well (: